### v1

Abstract
Generation of squeezed light is usually implemented in nonlinear or materials. Semiconductor lasers and optical amplifiers (SOAs) also offer nonlinearities but they differ from passive elements in that they add amplified spontaneous emission noise (ASE). In a semiconductor laser, squeezing to below the shot noise limit has been demonstrated. An SOA contains no cavity and it adds significant noise. Gain saturation can lead, in principle, to squeezing of the photon number quadrature to below the shot noise level but often the noise is reduced only to below the ASE level of a linear amplifier. At the same time, the noise in the phase quadrature increases according to the Heisenberg uncertainty principle. Short resonant pulses interacting with a quantum dot SOA induce coherent effects such as Rabi oscillations. Here, we demonstrate, for the first time, that Rabi oscillations cause cyclical noise squeezing which varies periodically with the excitation pulse area. The noise in the present experiments does not reach the quantum limit so we term this condition quasi-squeezing. It occurs during the portions of the Rabi cycle when the quantum dots provide gain and repeats with every fourfold increase of the pulse excitation energy which amounts to a 2 increase in pulse area. In all other cases, the noise exhibits the properties of a coherent state.
keywords:
squeezing, quantum dot, Rabi oscillations, semiconductor optical amplifier1 Introduction
Squeezed light is a non classical state of an optical signal in which the uncertainty (noise) in one quadrature (for example the photon number X) describing the field is reduced at the expense of the other quadrature (the phase P), while obeying the Heisenberg uncertainty [1, 2, 3]. Squeezed optical signals allow for improved fiber communication [4, 5, 6], and are imperative for precise measurements in quantum sensing and metrology [7], as well as for continuous variable quantum communication [8, 9]. Generation of squeezed light requires a strong nonlinearity which is most often implemented in passive or materials [10, 11, 12, 13, 14].
Strong nonlinearities are also available in optically active media, such as lasers and semiconductor optical amplifiers (SOAs). In the context of light squeezing, active media differ from passive materials in that they add spontaneous emission noise. The nonlinearity in an active device, and its consequent ability to squeeze the noise, stems from saturation which is determined, in turn, by the dynamical properties. In solid state and doped fiber devices, dynamics are slow and saturation is governed by average powers. In contrast, the dynamics of semiconductor devices are fast, and this has a major impact on the noise properties under saturated conditions [15, 16, 17].
Squeezing in a DC driven semiconductor laser to below the quantum limit was demonstrated by Y. Yamamoto’s group several decades ago [18, 19]. Similar squeezing levels were achieved in A. Yariv’s group by applying a weak dispersive feedback to the laser [20]. A model for squeezing in semiconductor lasers was reported in [21]. The noise of a saturated SOA was shown both theoretically and experimentally to exhibit a spectral hole in the vicinity of the saturating signal [22, 15, 23]. Clear narrowing of the the statistical distribution of the power spectral density in a saturated SOA was demonstrated experimentally for continuous waves (CW) [24] as well as pulse [25] excitations, and confirmed by a simulation [26].
Short pulse propagation in an SOA induces several nonlinear coherent interactions [27]. Once the pulses are sufficiently intense and their duration is short in comparison to the semiconductor coherence time (which is roughly 1 ps at room temperature [28]), they induce coherent interactions such as Rabi oscillations, which were demonstrated experimentally mainly in quantum-dot (QD) SOAs [27, 29, 30, 31], but also in a quantum cascade laser [32], and were modeled for quantum well SOAs [33]. This means that during the pulse, the SOA oscillates between gain and absorption. This leaves, in turn, a direct imprint on the temporal profiles of the amplitude and phase of the SOA output pulse. The carrier density in the SOA also oscillates and this affects the amplified spontaneous emission noise accompanying the pulse.
We demonstrate here, for the first time, a direct interplay between Rabi oscillations and noise squeezing in a QD SOA operating at 1550 nm and at room temperature. In the present experiments, gain saturation does not squeeze the noise below the shot noise limit but rather reduces it, in the photon number quadrature, to below its value under linear gain conditions. We term this condition quasi-squeezing. At the same time, the noise in the other quadrature is correspondingly increased.
To induce Rabi oscillations, we excite the QD SOA with 100 fs pulses and characterize the noise of the resulting output pulses using homodyne detection, where a replica of the excitation pulse acts as the local oscillator (LO). Noise squeezing is found to occur only under conditions of gain saturation and it is cyclical with the input pulse excitation energy. Namely, squeezed noise is observed when the excitation pulse energies increase by factors of four which amounts to a 2 increase in pulse area. For other pulse excitation energies, the QD SOA output exhibits a coherent state. This periodic behavior extends over a very wide range of excitation pulse energies.
2 Results and Discussion
The experimental setup for homodyne characterization of ultra short pulses following coherent interaction with a QD SOA is shown in Fig. 1. A 100 femtosecond laser pulse is split with one part serving as the LO and the second (signal) is coupled to a QD optical amplifier. The signal pulse interacts with the QD gain medium, and the transmitted field is recombined with the LO, whose delay is controlled via an optical delay line. The resulting interference is detected by a balanced photodetector, and the difference signal is recorded using an oscilloscope. Scanning the LO phase over 2 radians for about one thousand times yields a histogram of the current uncertainty. The histogram is transformed into a Wigner function using the Inverse Radon Transformation [34]. The insets in Fig. 1 show respectively the Wigner functions of the amplified spontaneous emission noise under linear conditions namely, with no input signal and of a typicsl quasi squeezed state.
The homodyne system serves also as a cross-correlator where the LO pulse samples the output pulses of the QD gain medium while the relative temporal delay between them is scanned. Such cross-correlations for three different excitation pulse energies are shown in the left column of Fig. 2. The cross-correlations reveal that the 100-fs input pulse broadens to about 1 ps and attains an amplitude profile resembling the results measured using cross-correlation frequency resolved optical gating (X-FROG) [29]. Homodyne measurements were performed with the LO sampling the signal at different temporal locations. The center column of Fig. 2 describes histograms of the normalized balanced homodyne detector output current as a function of the LO phase, measured at the trailing peak (dashed line at 0 ps). The corresponding extracted Wigner function are depicted in the right column, revealing a clear dependence on the input pulse energy. At 6.3 and 25 pJ, the Wigner function is that of a coherent state. However, at 12.5 pJ it exhibits clear quasi-squeezing. At any other temporal position, the Wigner functions for all input pulse energies exhibit coherent states namely, they are not affected by by the QD dynamics stemming from the Rabi oscillations.
The dependence on pulse excitation energy was examined in more detail as described in Fig. 3. The upper left panel shows, for reference, the Wigner function of the amplified spontaneous emission (ASE) noise obtained with no input signal. Quasi-squeezing is found to be cyclical with a periodicity of a fourfold increase in energy, from 3.96 pJ to 15.8 pJ corresponding to a pulse area change of 2. The two quasi squeezed Wigner functions at 3.96 pJ and 15.8 pJ exhibit a photon number uncertainty which is smaller than that of the ASE as seen by the dashed circles. Note that the absolute values of the pulse area cannot be determined precisely, since a QD SOA is an inhomogeneously broadened, distributed gain medium. The pulse area is defined as where is the dipole matrix element and is the envelope of the electric field at position . According to the McCall–Hahn area theorem [35], which was originally derived for resonant two-level systems under the assumptions of negligible dephasing, no inhomogeneous broadening, and no nonlinear gain dynamics, the pulse area in an amplifying medium evolves as where is the gain coefficient. In a QD amplifier, these assumptions are only qualitatively satisfied: the ensemble exhibits inhomogeneous broadening, nonlinearity due to gain saturation, and complex carrier dynamics beyond the two-level approximation. While the exact initial pulse area cannot be determined accurately, changes in pulse area of are nevertheless exact. Quasi-squeezing occurs only when the system is in the gain regime during the Rabi cycle, while in every other case, the amplifier output is a coherent state whose Wigner function size is similar to that of the amplified spontaneous emission noise.
The cyclical nature of the quasi squeezing is supported by a simple phenomenological calculations which analyzes the spontaneous emission of an ideal two level system driven by a short excitation pulse , whose area is defined as . The pulse induces Rabi oscillations at a frequency , where is the transition dipole moment and is the reduced Planck constant. In the rotating frame of the excitation pulse, the instantaneous excited-state population evolves as:
Spontaneous emission noise is incorporated by adding Gaussian noise whose variance is proportional to the excited-state population. The output electric field becomes:
Both noise terms, and take on the form of a normal distribution
where is the spontaneous emission rate and the two noise terms are related to each other via the Henry’s linewidth enhancement parameter [36].
To estimate the resulting quantum state, we compute the average field amplitude and phase from the noisy output: . This complex amplitude is scaled and used to construct a coherent density matrix . The Wigner function is then calculated directly from this density matrix using the standard QuTiP routine: , where denotes the Wigner function of state .
Simulated Wigner functions are shown in Fig. 4 for input pulse areas ranging from to 5. The periodical nature of quasi-squeezing is clearly observed for pulse areas that differ by 2, consistent with the measured data.
3 Conclussion
In this work, we demonstrated that a resonant ultra-short pulse excitation of a QD SOA induces cyclical quasi-squeezing, driven by coherent Rabi oscillations. Specifically, we observed that the output photon number quadrature noise is periodically reduced below the ASE level of a linear amplifier, while the conjugate phase quadrature noise increases correspondingly. This quasi-squeezing effect repeats with every fourfold increase in the excitation pulse energy, equivalent to a 2 increment in pulse area.
The underlying mechanism arises from the periodic evolution between gain and absorption in the QD SOA during Rabi oscillations. The noise is squeezed only during portions of the Rabi cycle when the QDs provide gain; otherwise, the output exhibits the noise properties of a coherent state. This behavior is in qualitative agreement with predictions based on the McCall–Hahn area theorem, despite the inhomogeneous broadening and complex carrier dynamics in QD SOAs.
While in the present experiments the noise does not reduce to below the quantum shot noise limit, it represents a clear manipulation of amplifier-added noise via coherent light–matter interactions. This demonstrates a fundamentally new regime of amplifier operation that could enable engineered noise profiles in active photonic devices. Potential applications include tailored noise suppression in optical amplifiers for quantum communication systems, or enhanced sensitivity in measurement schemes where reduced amplitude noise, is crucial.
Future research should explore extending this effect to other material systems, for example perovskite based amplifiers, and investigate whether cavity enhancement or tailored feedback could ensure squeezing below the shot noise limit. Additionally, incorporating pulse shaping techniques to precisely control the temporal dynamics of the Rabi oscillations [37] may open new avenues for deterministic quantum noise engineering on chip-scale photonic platforms.
4 Methods
4.1 Experimental setup
The experimental setup consists of a femtosecond fiber-based laser (Toptica FemtoFiber pro) operating at a center wavelength of 1550 nm, delivering pulses with a duration of approximately 100 fs. The optical path length was controlled using a motorized optical delay line (Physik Instrumente, model V-508, step resolution of 2 nm) to achieve precise temporal scanning. The output beams were combined and directed to a balanced photodetector (Femto, model HBPR-100M-60K-IN-FST) for detection of the interference signal. The resulting photocurrent was recorded using an oscilloscope (Keysight, model EXR404A, with 4 GHz bandwidth) for subsequent analysis. The delay line was moving periodically at rate of 5 Hz and an amplitude of 2 m and correspondingly produced a position signal that was sampled by the oscilloscope in sync with the signal from the balanced detector.
4.2 QD SOA fabrication
The active region of the gain medium we used comprised six layers of high-density cm-2 InAs QDs grown by molecular beam epitaxy in the Stranski - Krastanow mode. The optical amplifier is fabricated as a 2 wide, 1.5 mm long ridge waveguide whose end facets were anti-reflection coated. The QDs exhibit a narrow photo-luminescence linewidth, measured at 10 K, of 17 meV for a single layer and 26 meV for the six-layer stack. This linewidth represents a record uniformity compared to all reported QDs in any material system. Details of the device fabrication and characterization are given in [38]. The emission power of the measured electro-luminescence spectra increases with applied bias but the spectral shape remains unchanged [39].
4.3 Wigner function analysis
Histograms of the current at the balanced detector output, measured with a real time oscilloscope, is constructed from one thousand, periodic scans of the delay line with an amplitude of 2m. Wigner functions are obtained from those histograms using the inverse Radon transform. This transform is intended for continuous waves (CW) or narrow-spectrum signal. The spectral width of the 100 fs excitation pulses is roughly 3% of the optical frequency and hence, the assumption of narrow band signal is valid.
Acknowledgements The reearch was partially supported by the Israel Science Foundation under grant number 460/21 and by the Hellen Diller Quantum Center at Technion. IK acknowledges financial support from the Ministry of Innovation, Science and Technology of Israel.